OpenStudy (anonymous):

Factor completely using the difference of squares along with factoring by grouping. 4t^3-16y^2-9t+36

OpenStudy (anonymous):

do you mean 4t^3-16t^2-9t+36?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

4t^3-16t^2-9t+36 (4t^3-16t^2) + (-9t+36) i grouped them for you; what can you factor form each one

OpenStudy (anonymous):

what is the common factor in 4t^3-16t^2?

OpenStudy (anonymous):

4t^2

OpenStudy (anonymous):

correct factor it what do you get? 4t^2 ( ____ - _____)

OpenStudy (anonymous):

4t^2(t-4)

OpenStudy (anonymous):

factor this too (-9t+36)

OpenStudy (anonymous):

-9(t+4) ?

OpenStudy (anonymous):

nope -9(t-4) now we have 4t^2(t-4) + (-9)(t-4) what is the common factor now?

OpenStudy (anonymous):

(t-4)

OpenStudy (anonymous):

ok factor it

OpenStudy (anonymous):

(t - 4)(4t^2-9) do you know how to factor (4t^2-9)

OpenStudy (anonymous):

(a^2 - b^2)=(a+b)(a-b) ((2*t)^2 - 3^2) = (2t - 3)(2t + 3) now (2t - 3)(2t + 3)(t - 4)

OpenStudy (anonymous):

got it?